OpenStudy (anonymous):

what is the solution to the following equation x^2+2x+7=0

OpenStudy (amistre64):

there are a few ways to approach a solution, which ones have you tried?

OpenStudy (amistre64):

completing the square is the least "upsetting" to me

OpenStudy (amistre64):

"x^2 +2x +1 " is a complete square since it compresses to (x+1)^2 so lets borrow what we need from that 7 to make this: x^2 + 2x +1 + 6 = 0 (x+1)^2 + 6 = 0 and algebra it into submission